# Limits to scalable evaluation at the frontier: LLM as Judge won’t beat twice the data

### v1

@@ C.6 Proposition 4
We first note that $\operatorname*{\mathrm{Var}}s(m)=b(m)(1-b(m))$ and $\operatorname*{\mathrm{Var}}\tilde{s}(m)=\operatorname*{\mathbb{E}}\tilde{s}(m)(1-\operatorname*{\mathbb{E}}\tilde{s}(m)),$ as both $s$ and $\tilde{s}$ are binary. Furthermore, $\operatorname*{\mathbb{E}}[s(m)\tilde{s}(m)]=b(m)p(m)$, such that $\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))=b(m)(p(m)-\operatorname*{\mathbb{E}}\tilde{s}(m))$. With this, it is easy to see that the formula for $\rho^{2}$ in the proposition statement is indeed correct. With this, we take the limit of $N\to\infty$ to maximize the variance reduction and plug in our parameterization for $s$ and $\tilde{s}$ to simplify
$1-\frac{\operatorname*{\mathrm{Var}}\hat{\theta}^{PP}_{\lambda^{*}}}{\operatorname*{\mathrm{Var}}\hat{\theta}}=\rho(s(m),\tilde{s}(m))^{2}$ $=\frac{b(m)}{(1-b(m))}\frac{(p(m)-\operatorname*{\mathbb{E}}\tilde{s}(m))^{2}}{\operatorname*{\mathbb{E}}\tilde{s}(m)(1-\operatorname*{\mathbb{E}}\tilde{s}(m))}\,.\qed$

### v2

@@ C.7 Proposition 4
We first note that $\operatorname*{\mathrm{Var}}s(m)=b(m)(1-b(m))$ and $\operatorname*{\mathrm{Var}}\tilde{s}(m)=\operatorname*{\mathbb{E}}\tilde{s}(m)(1-\operatorname*{\mathbb{E}}\tilde{s}(m)),$ as both $s$ and $\tilde{s}$ are binary. Furthermore, $\operatorname*{\mathbb{E}}[s(m)\tilde{s}(m)]=b(m)p(m)$, such that $\operatorname*{\mathrm{Cov}}(s(m),\tilde{s}(m))=b(m)(p(m)-\operatorname*{\mathbb{E}}\tilde{s}(m))$. With this, it is easy to see that the formula for $\rho^{2}$ in the proposition statement is indeed correct.
We then note, that for fixed values of $n$ and $N$, the relative variance reduction of $\hat{\theta}^{PP}_{\lambda^{*}}$ compared to $\hat{\theta}^{GT}$ is determined by the squared Pearson correlation coefficient $\rho$ between the ground truth labels $s$ and model judgements $\tilde{s}$:
Using the definition of $\tau$, this yields
$\tau(\hat{\theta}^{PP}_{\lambda^{*}})=\frac{1}{1-\frac{1}{1+\frac{n}{N}}\rho(s(m),\tilde{s}(m))^{2}}.$
Finally, we take the limit of $N\to\infty$ for an upper bound. ∎
C.8 Proposition 8
Proof.
We drop the parameters’ dependence on $m$ and set $q=1-b$ and $p=b$. Then $\operatorname*{\mathrm{JB}}=b(1-b)-(1-b)b=0.$ At the same time, as there is no judge bias, $\operatorname*{\mathbb{E}}\tilde{s}(m)=b=p$, such that $\rho(s(m),\tilde{s}(m))^{2}=0$. Lastly $\operatorname*{\mathrm{AG}}=b^{2}+(1-b)^{2}$, which can attain any value between $0.5$ and $1$. ∎